Given P(E or F) = 0.89, P(F) = 0.44, and P(E and F) = 0.04, what is P(E)?
Answer #1

Solution:

P( E or F ) = P(E) + P(F) - P(E and F)

0.89 = P(E) + 0.44 - 0.04

P(E) = 0.89 - 0.40

P(E) = 0.49